# Digital Zero-Noise Extrapolation with Quantum Circuit Unoptimization

### v2

@@ Abstract
Quantum circuit unoptimization is an algorithm that transforms a quantum circuit into a different circuit that uses more gate operations while maintaining the same unitary transformation. We demonstrate that this method can implement digital zero-noise extrapolation (ZNE), a quantum error mitigation technique. By employing quantum circuit unoptimization as a form of circuit folding, noise can be systematically amplified. The key advantages of this approach are twofold. First, its ability to generate an exponentially increasing number of distinct circuit variants as the noise level is amplified, which allows noise averaging over many circuit instances with slightly different circuit structure which mitigates the effect of biased error propagation because of the significantly altered circuit structure from quantum circuit unoptimization, or highly biased local noise on a quantum processor. Second, quantum circuit unoptimization by design resists circuit simplification back to the original unmodified circuit, making it plausible to use ZNE in contexts where circuit compiler optimization is applied server-side. We evaluate the effectiveness of quantum circuit unoptimization as a noise-scaling method for ZNE in two test cases using depolarizing noise numerical simulations: random quantum volume circuits, where the observable is the heavy output probability, and QAOA circuits for the (unweighted) maximum cut problem on random 3-regular graphs, where the observable is the cut value. We show that using quantum circuit unoptimization to perform ZNE can approximately recover signal from noisy quantum simulations.
@@ 1 Introduction
Numerous quantum error mitigation algorithms have been developed, including probabilistic error cancellation (PEC) [26, 27, 28], Clifford data regression (CDR) [29], and machine learning-based approaches [6, 30, 31, 32, 33]. This study focuses on zero-noise extrapolation (ZNE) [7, 12, 14, 5, 26, 3], a quantum error mitigation technique that amplifies the noise in a quantum computation by a specified factor. The resulting measurements contain a controlled amount of noise or error. An observable quantity is extracted from these measurements, and an extrapolation is performed to estimate the zero-noise limit. ZNE in comparison to many other quantum error mitigation algorithms has a lower computational and sampling overhead, making it more accessible to use on current quantum computers. To summarize, ZNE is comprised of the following components:
@@
Generally, there are two methods for scaling the noise. The first is digital, meaning in the circuit representation of the unitary that is being simulated—typically, this comes in the form of an increased number of circuit instructions that are still implementing an identical unitary (under noiseless conditions). The second is by stretching the pulse-level waveforms on the hardware [5, 34] that are implementing the circuit-level instructions—in particular by still implementing the same operations, but over longer timescales, which thus creates more noise due to the limited qubit coherence times and an accumulation of errors over time.
In this study, we demonstrate that digital zero-noise extrapolation, specifically, the noise amplification portion of the algorithm, can be implemented using quantum circuit unoptimization, which was introduced in [35] to test the effectiveness of quantum circuit compiler optimization. Quantum circuit unoptimization has several valuable properties compared to the existing digital noise amplification methods.
First, quantum circuit unoptimization enables the generation of an exponentially growing number of potential circuit variants as the circuit noise is amplified. Circuit variant here means a circuit with a different set of elementary quantum gates, but still implements exactly the same unitary operation. In contrast, other methods, such as global unitary circuit folding, generate exactly one circuit per noise scale factor. This exponential growth in possible circuit variants can be particularly advantageous when the noise profile of the quantum hardware is highly variable or non-uniform. By averaging over multiple circuit variants—each representing a slightly different implementation of the same logical unitary operation—one can mitigate the impact of localized noise effects. This circuit variant averaging also mitigates noise biasing that can occur from the inherent differences in quantum circuit structure from the unoptimization routine. This approach reduces bias in the measured observables, as the diverse noise profiles, or specifically physical implementations, of the circuit effectively “sample” different noise environments, leading to more robust results.
@@
Third, quantum circuit unoptimization allows for fractional noise scale factors. While this is not a property unique to quantum circuit unoptimization, it is a significant advantage for ZNE. Fractional noise scaling enables finer control over the noise amplification process, particularly for scale factors close to $1.0$. This capability is especially valuable in highly noisy computation scenarios where slight deviations from the original circuit may already introduce significant noise. Unoptimization facilitates more accurate extrapolation to the zero-noise limit by enabling precise control over noise scaling.
@@ 2.1 Quantum Circuit Unoptimization
For the insert step, a random two-qubit unitary $A$ and its Hermitian conjugate $A^{{\dagger}}$ are introduced into the circuit. These gates are inserted between two existing two-qubit gates in the circuit, denoted as $B_{1}$ and $B_{2}$. The insertion of $A$ and $A^{{\dagger}}$ does not alter the logical functionality of the circuit, as the combination $A^{{\dagger}}A=I$. The selection of $B_{1}$ and $B_{2}$ can follow different strategies, such as choosing gates that share a common qubit (the concatenated strategy) or selecting them randomly (the random strategy). The choice of strategy affects the unoptimization procedure’s characteristics and its impact on circuit structure. The concatenated strategy, as we implement it, is specifically a deterministic two 2-qubit gate selection procedure, where we choose the first pair of two-qubit gates that share a common qubit. This means specifically that for the concatenated strategy, there is no randomness in the selection of this pair of $2$-qubit gates. However, one could certainly utilize random selection of pairs of two qubit gates that share a common qubit.
@@
representing the transformation of $A^{{\dagger}}$ under the influence of $B_{1}$. This step introduces additional complexity in the circuit by altering the placement and nature of the gates while maintaining unitary equivalence.
Now containing additional gates from the first two steps, the circuit undergoes a decomposition and gate synthesis process. These last two steps can be thought of as effectively grouped together because this is really just a circuit compilation routine to compile down to some quantum instruction gateset. Critically, these steps are expressing the unoptimized circuit, which is now comprised of Multi-qubit unitary gates, including $\widetilde{A^{{\dagger}}}$, are broken down into elementary gates, such as single-qubit rotations and CNOT gates. This decomposition is performed using Qiskit [36], specifically a single decomposition pass. The final step optimizes the decomposed circuit using a synthesis procedure, in this case we implement this step in the form of the Qiskit transpiler [36], with optimization level $3$ (which is the maximum), targeting a particular gateset. If this circuit is being implemented on a specific quantum computing architecture, then the native gateset of that quantum computer would be used as the target gateset when performing these decomposition and synthesis steps.
@@
Figure 1: Depiction of quantum circuit unoptimization applied to ZNE. Panel a) shows a small, but representative, example of quantum circuit unoptimization being used. The example is a 4-qubit fully connected graph state circuit which is the input for a single round of the circuit unoptimization algorithm, using the concatenated pair of two qubit gates selection strategy (the common qubit is $q_{0}$). Unoptimization consists of gate insert, swap, decompose, and synthesis operations. The right-most circuit shows the final state of the original circuit after a single round of unoptimization has been applied to it. The input and resulting circuits are unitarily equivalent to one another, which remains true even after repeated rounds of unoptimization. Recursive application of quantum circuit unoptimization results in a growing combinatorial explosion of possible circuit (equivalent) variants with different circuit structure as more iterations are applied. This large number of potential circuit variants is due to random choice of two 2-qubit gates during the Insert step (this is only true for the random strategy), and because the two qubit gate $A$ in the Insert step is chosen as a random unitary. Measurement gates are not shown for brevity in the circuit renderings, but in this case the state of all 4 qubits at the end of the circuit are measured. Additional information about the gates used in the Qiskit-generated circuit diagrams are provided in the Appendix A. Panel b) shows a conceptual diagram of using quantum circuit unoptimization to amplify noise and then carry out zero noise extrapolation, given an arbitrary input circuit $C$ consisting of single and two-qubit gates. For some choice of integers $i_{1}<i_{2}<\cdots<i_{n}$, we apply $i_{k}$ iterations of the recipe to the circuit $C$, denoted as $R_{i_{k}}(C)$ for some $k\in\{1,\ldots,n\}$. For each resulting circuit from $R_{i_{k}}(C)$, we obtain a corresponding noise-scale factor as a function of the resulting circuit’s depth and the original circuit’s depth. The greater the number of recipe iterations corresponds to a circuit with increasing depth (and hence increased noise). We compute the expectation value of each noise-scale factor and extrapolate to the zero-noise limit via some extrapolation method (e.g., linear, Richardson, exponential, etc.).
@@
The unoptimization recipe can be applied iteratively to a quantum circuit with each application subsequently increasing the circuit’s depth, increasing the total number of instructions, and altering the overall circuit structure. Let $R(C)$ represent the application of the unoptimization recipe to a circuit $C$. After $i$ iterations, the resulting circuit can be expressed recursively as
@@ 2.2 Test Cases
QAOA aims to approximate ground states of diagonal (classical) cost Hamiltonians corresponding to combinatorial optimization problems. However, for QAOA to perform well, a sufficiently large depth $p$ is required, and the QAOA angles must be carefully chosen. These parameters are typically optimized variationally, or designed to approximate adiabatic quantum evolution [57]. In this study, we use the fixed-angle QAOA parameters proposed in [58], which are known to perform well on random 3-regular Max-Cut instances due to the parameter concentration property of the QAOA control parameters which have been seen in several problem classes [59, 60, 61]. The observable of interest in our ZNE simulations is the expected graph cut value given by the Hamiltonian,
@@ 2.3 Numerical Simulations
We present several numerical simulations to analyze the noise characteristics of using quantum circuit unoptimization for ZNE. All numerical simulations are performed using depolarizing noise (error) models using the Qiskit quantum programming library [36] (version number 1.3.2). Specifically, a depolarizing error rate of $0.001$ is applied to all single and two-qubit gates and a finite shot count of $1\mathrm{e}{-6}$ measurements are used for each circuit and noise scale factor execution. The quantum circuit unoptimization noise scale factor (eq. (3)) is computed after applying the Qiskit transpiler optimization pass to each unoptimization iteration. The unoptimization iteration is applied to the original circuit after applying the Qiskit transpiler optimization pass. optimization_level $3$ was used for all Qiskit optimization passes. All circuits are transpiled to the basis gates of cx (CNOT) and U3 – reported single and two-qubit gate counts are based on only these two gates.
The extrapolation functions we test are linear and quadratic functions. The coefficients for these extrapolations are fit using least squares curve fitting. The quadratic function is $ax^{2}+bx+c$ where $a,b$ and $c$ are tunable coefficients. The linear function is $ax+b$ where $a$ and $b$ are tunable coefficients.
@@ 3 Results
Figure 3: Zero-noise extrapolation using linear and quadratic regression for a Quantum Volume circuit with $10$ qubits using quantum circuit unoptimization gate selection method random (right) and concatenated (left). There were $35$ circuit unoptimization iterations used to amplify the noise and each step’s estimated observable is plotted as a blue point in each sub-figure. Each sub-plot shows the heavy output probability (y-axis) as a function of the noise scale factor $\lambda$. The red marker denotes the noiseless observable quantity, and the two yellow markers denote the quantum circuit unoptimization ZNE estimates of that observable. Note that each sub-plot used a single run of recursive quantum circuit unoptimization – meaning that the unoptimized circuit built on itself as the number of steps increased, as opposed to independently choosing a new unoptimization path (with a new random seed) for each noise level.
@@
Figure 4: Zero-noise extrapolation using linear and quadratic regression for a fixed-angle $p=2$ QAOA circuit applied to Max-Cut on a random 3-regular graph with $12$ qubits using quantum circuit unoptimization gate selection method random (right) and concatenated (left). $35$ quantum circuit unoptimization iterations were used to amplify the noise (the x-axis shows the computed $\lambda$ quantity for each one of these iterations), and each step’s estimated observable is plotted in each sub-figure. Note that each sub-plot used a single run of recursive quantum circuit unoptimization – meaning that the unoptimized circuit built on itself as the number of steps increased, as opposed to independently choosing a new unoptimization path (with a new random seed) for each noise level.
Figure 5: Averaged circuit variant zero-noise extrapolation numerical simulation tests. For each sub-plot, quantum circuit unoptimization is applied using a total of $30$ independent recursive sequences (each from $1$ to $35$ iterations). This is unlike Figures 3 and 4, where only a single unoptimization sequence is used. The underlying quantum circuits (a fixed quantum volume circuit in the left column and a fixed Max-Cut QAOA circuit in the right-hand column), and observables, are the same as in Figures 3 and 4. Each separate quantum circuit unoptimization run can result in slightly different noise scale factors (eq. (3)). The random selection strategy applied to the $p=2$ QAOA circuit has the highest observable variability. Still, in all four instances, the zero-noise extrapolation can recover a good signal from the noisy quantum circuit simulation. Notice that the variance seen in the individual datapoints for the concatenated strategy plots is purely due to the random two qubit unitary selection during the insert stage of unoptimization.
Figure 3 shows quantum circuit unoptimization ZNE applied to a random quantum volume circuit. These results show that the noise can be amplified predictably and thus the noiseless signal can be recovered reasonably well.
@@
Combined, these results show that the concatenated strategy is better for performing controlled (predictable) noise amplification using quantum circuit unoptimization, but also results in higher error rates than the random gate selection option as seen in Figure 6, due to the lack of random $2$-qubit gate selection that can result in better noise averaging. Therefore, the best strategy that results in the overall lowest error rate is to use the random gate selection strategy, and to average over many independent recursions of the quantum circuit unoptimization, and then to use the extrapolation that gives the best curve fit to the observable that is being measured.

### v3

@@ Abstract
Quantum circuit unoptimization is an algorithm that transforms a quantum circuit into a different circuit that uses more gate operations while maintaining the same unitary transformation. We demonstrate that this method can implement digital zero-noise extrapolation (ZNE), a quantum error mitigation technique. By employing quantum circuit unoptimization as a form of circuit folding, noise can be systematically amplified. The key advantages of this approach are twofold. First, its ability to generate an exponentially increasing number of distinct circuit variants as the noise level is amplified, which allows noise averaging over many circuit variants with slightly different circuit structure. Averaging over these variants can mitigate the effect of biased error propagation due to the significantly altered circuit structure from quantum circuit unoptimization, or biased noise sources on a quantum processor. Second, quantum circuit unoptimization by design resists circuit simplification back to the original unmodified circuit, making it plausible to use ZNE in contexts where circuit compiler optimization is applied server-side. We evaluate the effectiveness of quantum circuit unoptimization as a noise-scaling method for ZNE in two test cases using depolarizing noise numerical simulations: random quantum volume circuits, where the observable is the heavy output probability, and QAOA circuits for the (unweighted) maximum cut problem on random 3-regular graphs, where the observable is the cut value. We show that using quantum circuit unoptimization to perform ZNE can approximately recover signal from noisy quantum simulations.
@@ 1 Introduction
Numerous quantum error mitigation algorithms have been developed, including probabilistic error cancellation (PEC) [26, 27, 28], Clifford data regression (CDR) [29], and machine learning-based approaches [6, 30, 31, 32, 33]. This study focuses on zero-noise extrapolation (ZNE) [7, 12, 14, 5, 26, 3], a quantum error mitigation technique that amplifies the noise in a quantum computation by a specified factor. The resulting measurements contain a controlled amount of noise or error. An observable quantity is extracted from these measurements, and an extrapolation is performed to estimate the zero-noise limit. ZNE, in comparison to many other quantum error mitigation algorithms, has a lower computational and sampling overhead, making it more accessible to use on current quantum computers. To summarize, ZNE is comprised of the following components:
@@
Generally, there are two methods for scaling the noise. The first is digital, meaning in the circuit representation of the unitary that is being simulated—typically, this comes in the form of an increased number of circuit instructions that are still implementing an identical unitary (under noiseless conditions). The second is by stretching the pulse-level waveforms on the hardware [5, 34] that are implementing the circuit-level instructions—in particular, by still implementing the same operations, but over longer timescales, which thus creates more noise due to the limited qubit coherence times and an accumulation of errors over time.
In this study, we demonstrate that digital zero-noise extrapolation, specifically, the noise amplification portion of ZNE, can be implemented using quantum circuit unoptimization, which was introduced in [35] to test the effectiveness of quantum circuit compiler optimization. Quantum circuit unoptimization has several valuable properties compared to existing digital noise amplification methods.
First, quantum circuit unoptimization enables the generation of an exponentially growing number of potential circuit variants as the circuit noise is amplified. Circuit variant here means a circuit with a different set of elementary quantum gates, but still implements exactly the same unitary operation. In contrast, other methods, such as global unitary circuit folding, generate exactly one circuit per noise scale factor. This exponential growth in possible circuit variants can be particularly advantageous when the noise profile of the quantum hardware is highly variable or non-uniform. By averaging over multiple circuit variants—each representing a slightly different implementation of the same logical unitary operation—one can mitigate the impact of localized noise effects. This circuit variant averaging also mitigates noise biasing that can occur from the inherent differences in quantum circuit structure from the unoptimization routine. This approach reduces bias in the measured observables, as the diverse noise profiles, or specifically physical implementations, of the circuit effectively “sample” different noise environments, leading to more robust statistics.
@@
Third, quantum circuit unoptimization allows for fractional noise scale factors. While this is not a property unique to quantum circuit unoptimization, it is a significant advantage for ZNE. Fractional noise scaling enables finer control over the noise amplification process, particularly for scale factors close to $1.0$. This capability is especially valuable in highly noisy computation scenarios where slight deviations from the original circuit may already introduce significant noise. Quantum circuit unoptimization facilitates accurate extrapolation to the zero-noise limit by enabling precise control over noise scaling.
@@ 2.1 Quantum Circuit Unoptimization
For the insert step, a random two-qubit unitary $A$ and its Hermitian conjugate $A^{{\dagger}}$ are introduced into the circuit. These gates are inserted between two existing two-qubit gates in the circuit, denoted as $B_{1}$ and $B_{2}$. The insertion of $A$ and $A^{{\dagger}}$ does not alter the logical functionality of the circuit, as the combination $A^{{\dagger}}A=I$. The selection of $B_{1}$ and $B_{2}$ can follow different strategies, such as choosing gates that share a common qubit (the concatenated strategy) or selecting them randomly (the random strategy). The choice of strategy affects the unoptimization procedure’s characteristics and its impact on circuit structure. The concatenated strategy, as we implement it, is specifically a deterministic two 2-qubit gate selection procedure, where we choose the first pair of two-qubit gates that share a common qubit. This means specifically that for the concatenated strategy, there is no randomness in the selection of this pair of $2$-qubit gates. However, one could certainly utilize random selection of pairs of two-qubit gates that share a common qubit.
@@
representing the transformation of $A^{{\dagger}}$ under the influence of $B_{1}$. This step introduces additional complexity in the circuit by altering the placement and the nature of the gates while maintaining unitary equivalence.
@@
Figure 1: Depiction of quantum circuit unoptimization applied to ZNE. Panel a) shows a small, but representative, example of quantum circuit unoptimization being used. The example is a 4-qubit fully connected graph state circuit which is the input for a single round of the circuit unoptimization algorithm, using the concatenated pair of two-qubit gates selection strategy (the common qubit is $q_{0}$). Unoptimization consists of gate insert, swap, decompose, and synthesis operations. The right-most circuit shows the final state of the original circuit after a single round of unoptimization has been applied to it. The input and resulting circuits are unitarily equivalent to one another, which remains true even after repeated rounds of unoptimization. Recursive application of quantum circuit unoptimization results in a growing combinatorial explosion of possible circuit (equivalent) variants with different circuit structure as more iterations are applied. This large number of potential circuit variants is due to random choice of two 2-qubit gates during the Insert step (this is only true for the random strategy), and because the two-qubit gate $A$ in the Insert step is chosen as a random unitary. Measurement gates are not shown for brevity in the circuit renderings, but in this case the state of all 4 qubits at the end of the circuit are measured. Additional information about the gates used in the Qiskit-generated circuit diagrams is provided in the Appendix A. Panel b) shows a conceptual diagram of using quantum circuit unoptimization to amplify noise and then carry out zero noise extrapolation, given an arbitrary input circuit $C$ consisting of single and two-qubit gates. For some choice of integers $i_{1}<i_{2}<\cdots<i_{n}$, we apply $i_{k}$ iterations of the recipe to the circuit $C$, denoted as $R_{i_{k}}(C)$ for some $k\in\{1,\ldots,n\}$. For each resulting circuit from $R_{i_{k}}(C)$, we obtain a corresponding noise-scale factor as a function of the resulting circuit’s depth and the original circuit’s depth. The greater the number of recipe iterations corresponds to a circuit with increasing depth (and hence increased noise). We compute the expectation value of each noise-scale factor and extrapolate to the zero-noise limit via some extrapolation method (e.g., linear, Richardson, exponential, etc.).
Now containing additional gates from the first two steps, the circuit undergoes a decomposition and gate synthesis process. These last two steps can be thought of as effectively grouped together because this is just a circuit compilation routine to compile down to some quantum instruction gateset. Critically, these steps are expressing the unoptimized circuit, which is now comprised of Multi-qubit unitary gates, including $\widetilde{A^{{\dagger}}}$, and are broken down into elementary gates, such as single-qubit rotations and CNOT gates. This decomposition is performed using Qiskit [36], specifically a single decomposition pass. The final step optimizes the decomposed circuit using a synthesis procedure, in this case, we implement this step in the form of the Qiskit transpiler [36], with optimization level $3$ (which is the maximum), targeting a particular gateset. If this circuit is being implemented on a specific quantum computing architecture, then the native gateset of that quantum computer would be used as the target gateset when performing these decomposition and synthesis steps.
@@
The unoptimization recipe can be applied iteratively to a quantum circuit with each application subsequently increasing the circuit’s depth, increasing the total number of instructions and altering the overall circuit structure. Let $R(C)$ represents the application of the unoptimization recipe to a circuit $C$. After $i$ iterations, the resulting circuit can be expressed recursively as
@@ 2.2 Test Cases
QAOA aims to approximate ground state(s) of diagonal (classical) cost Hamiltonians corresponding to a combinatorial optimization problem. However, for QAOA to perform well, a sufficiently large depth $p$ is required, and the QAOA angles must be optimized. These parameters are typically optimized variationally, or designed to approximate adiabatic quantum evolution [57]. In this study, we use the fixed-angle QAOA parameters proposed in [58], which are known to perform well on random 3-regular Max-Cut instances due to the parameter concentration property of the QAOA control parameters, which have been seen in numerous problem classes [59, 60, 61]. The observable of interest in our ZNE simulations is the expected graph cut value given by the Hamiltonian,
@@ 2.3 Numerical Simulations
We present several numerical simulations to analyze the noise characteristics of using quantum circuit unoptimization for ZNE. All numerical simulations are performed using depolarizing noise (error) models using the Qiskit quantum programming library [36] (version number 1.3.2). Specifically, a depolarizing error rate of $0.001$ is applied to all single and two-qubit gates, and a finite shot count of $1\mathrm{e}{-6}$ measurements are used for each circuit and noise scale factor execution. The quantum circuit unoptimization noise scale factor (eq. (3)) is computed after applying the Qiskit transpiler optimization pass to each unoptimization iteration. The unoptimization iteration is applied to the original circuit after applying the Qiskit transpiler optimization pass. optimization_level $3$ was used for all Qiskit optimization passes. All circuits are transpiled to the basis gates of cx (CNOT) and U3 – reported single and two-qubit gate counts are based on only these two gates.
The extrapolation functions we test are linear and quadratic functions. The coefficients for these extrapolations are fit using least squares curve fitting. The quadratic function is $ax^{2}+bx+c$ where $a,b$, and $c$ are tunable coefficients. The linear function is $ax+b$, where $a$ and $b$ are tunable coefficients.
@@ 3 Results
Figure 3: Zero-noise extrapolation using linear and quadratic regression for a Quantum Volume circuit with $10$ qubits using quantum circuit unoptimization gate selection method, random (right) and concatenated (left). There were $35$ circuit unoptimization iterations used to amplify the noise and each step’s estimated observable is plotted as a blue point in each sub-figure. Each sub-plot shows the heavy output probability (y-axis) as a function of the noise scale factor $\lambda$. The red marker denotes the noiseless observable quantity, and the two yellow markers denote the quantum circuit unoptimization ZNE estimates of that observable. Note that each sub-plot used a single run of recursive quantum circuit unoptimization – meaning that the unoptimized circuit builds on itself as the number of steps increased, as opposed to independently choosing a new unoptimization path (with a new random seed) for each noise level.
Figure 4: Zero-noise extrapolation using linear and quadratic regression for a fixed-angle $p=2$ QAOA circuit applied to Max-Cut on a random 3-regular graph with $12$ qubits using quantum circuit unoptimization gate selection method random (right) and concatenated (left). $35$ quantum circuit unoptimization iterations were used to amplify the noise (the x-axis shows the computed $\lambda$ quantity for each one of these iterations), and each step’s estimated observable is plotted in each sub-figure. Note that each sub-plot used a single run of recursive quantum circuit unoptimization – meaning that the unoptimized circuit builds on itself as the number of steps increased, as opposed to independently choosing a new unoptimization path (with a new random seed) for each noise level.
@@
Figure 3 shows quantum circuit unoptimization ZNE applied to a random quantum volume circuit. These results show that the noise can be amplified predictably, and thus the noiseless signal can be recovered reasonably well.
@@
Figure 5: Averaged circuit variant zero-noise extrapolation numerical simulation tests. For each sub-plot, quantum circuit unoptimization is applied using a total of $30$ independent recursive sequences (each from $1$ to $35$ iterations). This is unlike Figures 3 and 4, where only a single unoptimization sequence is used. The underlying quantum circuits (a fixed quantum volume circuit in the left column and a fixed Max-Cut QAOA circuit in the right-hand column), and observables, are the same as in Figures 3 and 4. Each separate quantum circuit unoptimization run can result in slightly different noise scale factors (eq. (3)). The random selection strategy applied to the $p=2$ QAOA circuit has the highest observable variability. Still, in all four instances, the zero-noise extrapolation can recover a good signal from the noisy quantum circuit simulation. Notice that the variance seen in the individual data points for the concatenated strategy plots is purely due to the random two-qubit unitary selection during the insert stage of unoptimization.
Combined, these results show that the concatenated strategy is better for performing controlled (predictable) noise amplification using quantum circuit unoptimization, but also results in higher error rates than the random gate selection option as seen in Figure 6, due to the lack of random $2$-qubit gate selection that can result in better noise averaging. Therefore, the best set of parameters that result in the overall lowest error rate is to use the random gate selection strategy, then average over many independent recursions of quantum circuit unoptimization, and then to find a low-order polynomial extrapolation that gives the best curve fit to the observable that is being measured.